OpenStudy (anonymous):

How do I make the system (y+x/2=8) (2y=-x+16) into slope form (y=mx+b)?

OpenStudy (sheraz12345):

Can u make ur question more clear? Cant understand that system thingy

OpenStudy (yttrium):

By equation those systems to y. Let's take the first example.

OpenStudy (anonymous):

Its solving each system by graphing.. so first i have to make the equations into slope form (y=mx+b).

OpenStudy (yttrium):

\[y + \frac{ x }{ 2 } =8 \] Therefore, \[y = -\frac{ 1 }{ 2 }x + 8\]

OpenStudy (anonymous):

How did it become -1/2x+8?

OpenStudy (owlcoffee):

well, let's make them individually ito tht form: for the first equation: \[y+\frac{ x }{ 2 }=8\] and sustracting x/2 on both sides: \[y=- \frac{ x }{ 2 }+8\] for the second: \[2y=-x+16\] dividing both sides by two: \[y=\frac{ -x+16 }{ 2 }\] simplifying: \[y=-\frac{ x }{ 2 }+8\]

OpenStudy (yttrium):

By transposin it or like. \[y + \frac{ 1 }{ 2 }x -\frac{ 1 }{ 2 }x = \frac{ 1 }{ 2 }x + 8\]

OpenStudy (yttrium):

Therefore \[y = \frac{ 1 }{ 2 }x + 8\]

OpenStudy (anonymous):

Oh ok I get it now. So I guess its the same for the second one. Thanks Guys !